OpenStudy (anonymous):

the length of a rectangle is 5 times its width. if the perimeter of the rectangle is 60ft, find its area. need help again!

OpenStudy (lgbasallote):

let l = length w = width perimeter = l + l + w + w = 60 it says length is 5 times its width so l is 5 times w l = 5w perimeter = 5w + 5w + w + w = 60 solve for w

OpenStudy (lgbasallote):

then tell me what you got

OpenStudy (anonymous):

i dont get it because dont you need to know what w is? thats where im confuded i dont know how to solve it..

OpenStudy (lgbasallote):

that's why you solve for w. to get the value of w. make sense?

OpenStudy (anonymous):

ok but why does 5 work into that?

OpenStudy (lgbasallote):

because perimeter is l + l + w + w right?

OpenStudy (anonymous):

ya

OpenStudy (lgbasallote):

and it was said that length is 5 times its width so l = 5w i just put in that value of l into l + l + w + w 5w + 5w + w + w = perimeter

OpenStudy (lgbasallote):

got it now?

OpenStudy (anonymous):

soooo its just 5x5?

OpenStudy (anonymous):

im guessing im wrong..

OpenStudy (lgbasallote):

why is there a multiplication?

OpenStudy (anonymous):

i just thought because it has to equal 60..

OpenStudy (lgbasallote):

5w + 5w + w + w = 60 10 w + 2w = 60 12w = 60 solve for w

OpenStudy (anonymous):

5?

OpenStudy (lgbasallote):

right w is 5

OpenStudy (anonymous):

because 12 divided by 60 is 5

OpenStudy (lgbasallote):

now solve for l

OpenStudy (anonymous):

what does the problem look like?

OpenStudy (anonymous):

solving for l i mean

OpenStudy (anonymous):

..... you there?

OpenStudy (lgbasallote):

lol sorry..too many notifications i didnt notice you replied

OpenStudy (lgbasallote):

remember that thing i said a while ago? length is 5 times its width l = 5w

OpenStudy (anonymous):

its fine haha ok then l is five?

OpenStudy (lgbasallote):

not exactly...what's your w?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

but if l=5w theres nothing to solve there

OpenStudy (lgbasallote):

substitute 5 into w

OpenStudy (anonymous):

so l equals 5 times 5?

OpenStudy (jiteshmeghwal9):

Let the width be 'x'. then given:l=5*x=5x perimeter of a rectangle=2(l+b) A/q, 2(5x+x)=60 10x+2x=60 12x=60 x=60/12=5 now plug the value of x in the l & w. width=x=5 length=5*5=25 area=l*b 25*5=125 Ans.

OpenStudy (jiteshmeghwal9):

@cdub353 gt it???????

OpenStudy (lgbasallote):

yes @cdub353

OpenStudy (lgbasallote):

so what would be l then?

OpenStudy (anonymous):

ok ok ok this is making more sense now thanks to both of you!

OpenStudy (jiteshmeghwal9):

are u asking to me @lgbasallote ??

OpenStudy (anonymous):

25 and then 25x5=125

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

now you got it ^_^

OpenStudy (anonymous):

thank you!

OpenStudy (jiteshmeghwal9):

yw:)

OpenStudy (anonymous):

sooo now i have a slope question

OpenStudy (lgbasallote):

lol =)) post it as a new question. youll get more help that way

OpenStudy (anonymous):

ok haha